# Probing Deep into Temporal Profile Makes the Infrared Small Target Detector Much Better

### v1

@@ 4.3.4 Analysis of the level structure
To explore the effect of the level of the DeepPro structure, we removed some levels shown in Fig. 9, and the comparison results are in Table VII. It can be observed that fewer levels result in higher false alarm rates and lower detection rates, while the multi-level structure significantly improves the detection performance. This is because, with the decrease in the spatial resolution of the features, the process of a target moving through a specified location becomes slow and obvious. Meanwhile, the interference of noise is reduced since most noise is spatially uncorrelated. Consequently, the signal of the target on the temporal profile becomes broader and more distinctive, making it easier for the model to probe the anomaly signals. Furthermore, the multi-level architecture enables the network to capture more comprehensive temporal profile information of the target regions, enriching the feature representation.
@@ 4.4 Robustness to noise intensities
As the intensity of the noise increases, the detection performances of all methods experience degradations. Under extreme noise conditions, Res-UNet+DTUM suffers an obvious increase in terms of $F_{a}$, whereas our DeepPro consistently maintains a low $F_{a}$. In all conditions, our DeepPro achieves significantly superior detection performance. That is because the temporal probe mechanism can learn more reliable and richer correlation features of different signals in the temporal profile. Besides, the characteristics of long-term temporal variation enable robust discrimination between varying levels of noise, clutter, and target signals. Therefore, our DeepPro is robust to different noise intensity conditions.
@@ 4.5.1 Comparison to the state-of-the-arts
Models $SNR\leq 3$ NUDT-MIRSDT NUDT-MIRSDT-HiNo Params GFLOPs FPS
@@ 4.7 Failure cases
Figure 17: Failure cases in the real world and their gray variations on the temporal profile. (m1) is a pedestrian obscured by trees. (m2), (m3), and (m4) are stationary pedestrians. Their signals on the temporal profile are different from the ordinary moving target signals as shown in Figs. 3, 4 and 5.
Miss detecting obscured targets. The target (m1) in Fig. 17 is obscured by trees, leading to an incomplete signal on the temporal profile. The target signal is so narrow as a spike. This target is missed, and the reason can be twofold. On the one hand, this may be due to the incomplete signal. On the other hand, it may be due to the lack of such samples in the training set.
@@ 5 Conclusion
In this paper, we consider the essential domain and information for precise, universal, robust and efficient infrared small target (IRST) detection. We observe the targets from the temporal profile. We reveal some overlooked but essential temporal profile information, that is, the global temporal saliency of the target and the correlation information between target signal and other signals. We built the first prediction attribution tool, and validated the robustness of the temporal profile information (long-term temporal information) by applying it to IRST detection networks. Based on our research, we remodel the IRST detection task as a one-dimensional signal anomaly detection task and propose an efficient and effective IRST detection network, deep temporal probe network (DeepPro), in which the calculations (i.e., multiplication and addition) are only in the time dimension. Numerous experiments on widely-used benchmarks were conducted to confirm the effectiveness of our methods. It is exciting that our DeepPro outperforms the state-of-the-art methods especially on dim targets and in complex scenarios with extremely few parameters and extremely high processing efficiency. Through this work, we reveal the paramount importance of temporal profile information.

### v2

@@ 4.3.4 Analysis of the level structure
To explore the effect of the level of the DeepPro structure, we removed some levels shown in Fig. 9, and the comparison results are in Table VII. It can be observed that fewer levels result in higher false alarm rates and lower detection rates, while the multi-level structure significantly improves the detection performance. This is because, with the decrease in the spatial resolution of the features, the process of a target moving through a specified location becomes slow and obvious. Meanwhile, the interference of noise is reduced since most noise is spatially uncorrelated. Consequently, the signal of the target in the temporal profile becomes broader and more distinctive, making it easier for the model to probe the anomaly signals. Furthermore, the multi-level architecture enables the network to capture more comprehensive temporal profile information of the target regions, enriching the feature representation.
@@ 4.4 Robustness to noise intensities
As the intensity of the noise increases, the detection performances of all methods experience degradations. Under extreme noise conditions, Res-UNet+DTUM suffers an obvious increase in terms of $F_{a}$, whereas our DeepPro consistently maintains a low $F_{a}$. In all conditions, our DeepPro achieves significantly superior detection performance. That is because, the temporal probe mechanism can learn more reliable and richer correlation features of different signals in the temporal profile. Besides, the characteristics of long-term temporal variation enable robust discrimination between varying levels of noise, clutter, and target signals. Therefore, our DeepPro is robust to different noise intensity conditions.
@@ 4.5.1 Comparison to the state-of-the-arts
Models $SNR\leq 3$ NUDT-MIRSDT NUDT-MIRSDT-HiNo #Params GFLOPs FPS
@@ 4.7 Failure cases
Figure 17: Failure cases in the real world and their gray variations in the temporal profile. (m1) is a pedestrian obscured by trees. (m2), (m3), and (m4) are stationary pedestrians. Their signals in the temporal profile are different from the ordinary moving target signals as shown in Figs. 3, 4 and 5.
Miss detecting obscured targets. The target (m1) in Fig. 17 is obscured by trees, leading to an incomplete signal in the temporal profile. The target signal is so narrow as a spike. This target is missed, and the reason can be twofold. On the one hand, this may be due to the incomplete signal. On the other hand, it may be due to the lack of such samples in the training set.
@@ 5 Conclusion
In this paper, we consider the essential domain and information for precise, robust, and efficient infrared small target (IRST) detection. We observe the targets from the temporal profile. We reveal some overlooked but essential temporal profile information, that is, the global temporal saliency of the target and the correlation information between target signal and other signals. We built the first prediction attribution tool, and validated the robustness of the temporal profile information (long-term temporal information) by applying it to IRST detection networks. Based on our research, we remodel the IRST detection task as a one-dimensional signal anomaly detection task and propose an efficient and effective IRST detection network, deep temporal probe network (DeepPro), in which the calculations (i.e., multiplication and addition) are only in the time dimension. Numerous experiments on widely-used benchmarks were conducted to confirm the effectiveness of our methods. It is exciting that our DeepPro outperforms the state-of-the-art methods especially on dim targets and in complex scenarios with extremely few parameters and extremely high processing efficiency. Through this work, we reveal the paramount importance of temporal profile information.